# Ringdown signatures of Kerr black holes immersed in a magnetic field

### v2

@@ 4.3 Ringdown analysis and discussion
Table 4: Table of values comparing the IMR results for Kerr221 with the EW1221 results for the redshifted final mass $M_{f}$, dimensionless final spin $\tilde{a}_{f}$ and magnetic field $\tilde{B}_{f}$. For $M_{f}$ and $\tilde{a}_{f}$ we report the median value along with the range of the symmetric $90\%$ credible interval [85]. For the magnetic field $\tilde{B}_{f}$, we report the upper limit of the $90\%$ credible interval.
@@
GW150914 $68.8^{+3.6}_{-3.1}$ $71.7^{+13.2}_{-12.5}$ $67.5^{+10.6}_{-9.4}$ $0.69^{+0.05}_{-0.04}$ $0.69^{+0.18}_{-0.36}$ $0.61^{+0.19}_{-0.27}$ $\leq 0.79$
GW190521 $256.6^{+36.6}_{-30.4}$ $284.0^{+40.4}_{-43.9}$ $289.13^{+63.3}_{-47.1}$ $0.71^{+0.05}_{-0.07}$ $0.78^{+0.10}_{-0.22}$ $0.79^{+0.13}_{-0.21}$ $\leq 0.64$
GW190521_074359 $88.1^{+4.3}_{-4.9}$ $86.4^{+14.1}_{-14.8}$ $87.03^{+22.5}_{-17.0}$ $0.72^{+0.05}_{-0.07}$ $0.67^{+0.17}_{-0.34}$ $0.67^{+0.22}_{-0.39}$ $\leq 0.76$
GW200224_222234 $90.3^{+6.4}_{-6.3}$ $88.6^{+15.5}_{-15.2}$ $86.7^{+24.6}_{-15.7}$ $0.73^{+0.06}_{-0.07}$ $0.60^{+0.23}_{-0.42}$ $0.56^{+0.32}_{-0.44}$ $\leq 0.80$
Figure 6: Posterior distribution for the remnant mass and spin resulting from the analysis of GW150914 for three different ringdown models and the results from the full LVK IMR model using IMRPhenomXPHM (obtained from the data release [86]).
Having described the essential tools of our methodology, we now shift to analyzing LVK data provided by the GW Open Science Center [87, 88]. Employing the described framework, and assuming that the remnant object to be a EW1 remnant described by scalar quasinormal modes, we conduct an analysis of the ringdown data for four binary black hole merger events. The events were chosen based on their high signal to noise and sufficiently long waveform for a ringdown-only analysis.
The results are shown in the corner plots of Fig. 5, which exhibit the $90\%$ credible interval of the two-dimensional posterior distributions on remnant redshifted mass, spin and magnetic field for the set of four selected events. The results of which are also tabulated in Table 4, where they are compared to results determined by LVK using the standard Kerr template. Finally, in Fig. 6, we show the two-dimensional posterior for the event GW150914 to compare how different models impact the extraction of the remnant spin $\tilde{a}_{f}$ and mass $M_{f}$. The chosen models for comparison include: scalar QNM for linearized Kerr at $\mathcal{O}(\tilde{a})$, tensor QNM for Kerr at $\mathcal{O}(\tilde{a}^{2})$, scalar QNM for linearized Ernst-Wild at a chosen value of $\tilde{B}_{f}=0.3$ and the full LVK inspiral-merger-ringdown (IMR) posterior determined using the IMRPhenomXPHM model [89].
The results indicate no evidence for a deviation from the Kerr hypothesis, and while the ensuing bounds are relatively weak, the current data already provides sensitivity to magnetic fields above $\tilde{B}\sim 0.3$. This is still very large in astrophysical terms, but indicates that the presence of a magnetosphere may provide one further factor complicating the extraction of black hole parameters from the merger. The exploratory nature of this analysis, apparent in the approximations A1 and A2, limits the scope of broader conclusions, but relaxing both assumptions is technically feasible, as discussed in the next section.
@@ 5 Concluding Remarks
Our results indicate consistency with the LVK analysis [90] using the Kerr template, and the level of sensitivity to the magnetic field indicates that it will impact the extraction of parameters at current levels of sensitivity only if relatively large in units of the maximal field $B_{M}\sim 1/M$. However, our analysis has relied critically on the approximations A1 and A2, and it is natural to ask about the potential impact of relaxing these assumptions. It is apparent that the perturbative analysis at ${\cal O}(\tilde{a})$ is a significant restriction given the current dataset, where the majority of events have relatively high spin, so that extending to ${\cal O}(\tilde{a}^{2})$ would be valuable. However, this may change with future observing runs (and eventually with the launch of space interferometers like LISA) potentially detecting events with smaller $\tilde{a}$, perhaps from black holes that have not merged before [91]. Separately, based on the Kerr examples, switching from scalar to tensor modes may also induce non-negligible shifts in the QNM frequencies, and of course a fully quantitative analysis would require the physical tensor QNMs. Beyond the assumptions A1 and A2 that directly impact the ringdown template, there is also the problem that currently there is no information about the pre-merger system for producing the EW magnetized remnant, and specifically how the inspiral and merger may differ from the Kerr template. Further information about the evolution of the magnetosphere during the inspiral would also be needed for a more quantitative analysis, to gain insight into which QNMs are more likely to be excited in the remnant, and to better model the start time $t_{0}$ for the ringdown analysis.
@@ Acknowledgements.
We are particularly grateful to Gregorio Carullo for assistance with the implementation of pyRing, and to Michel Lefebvre for many valuable discussions regarding aspects of this work and for carefully reading the manuscript. We would also like to thank Emanuele Berti for illuminating discussions at the Testing Gravity 2023 conference, Frans Pretorius for providing additional insight at the Werner Israel Memorial Symposium, and Michalis Agathos for helpful comments at the poster session at the Kavli-Villum Summer School on Gravitational Waves. This work was supported in part by NSERC, Canada, and has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/), a service of the LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. Software: This work makes use of the following python packages: pyRing, corner, cpnest, GWpy, lalsuite, matplotlib, numpy, scipy, and seaborn [66, 84, 92, 93, 94, 95, 96, 97].
@@ Appendix A Ernst-Wild asymptotics at ${\cal O}(\tilde{B}^{2})$
and the Klein Gordon equation becomes [37],
@@
The effective potential for the Schwarzschild black hole on the first line is only corrected by a constant at large radius at ${\cal O}(\tilde{B}^{2})$. The shift is exhibited in Fig. 7, with the shaded region showing the domain beyond the Melvin radius where the asymptotics would be modified more substantially at ${\cal O}(\tilde{B}^{4})$. The restriction to ${\cal O}(\tilde{B}^{2})$ thus restores the expectation within the Kerr spacetime that the QNM spectrum is primarily affected by the behaviour of the potential near the horizon and photon sphere of the black hole [98, 99, 63, 100, 37].

### v3

@@ 4.3 Ringdown analysis and discussion
Table 4: Table of values comparing the IMR results for Kerr221 [KAGRA:2023pio, LIGOScientific:2019lzm, LIGOScientific:2020tif, LIGOScientific:2021sio] with the EW1221 results for the redshifted final mass $M_{f}$, dimensionless final spin $\tilde{a}_{f}$ and dimensionless magnetic field $\tilde{B}_{f}$, where the latter can be restored to physical units using Eq. 5. For $M_{f}$ and $\tilde{a}_{f}$, we report the median value along with the range of the $90\%$ credible interval [LIGOScientific:2013yzb]. For the magnetic field $\tilde{B}_{f}$, we report the upper limit of the $90\%$ credible interval.
@@
IMR $\mathrm{Kerr_{221}}$ $\mathrm{EW1_{221}}$ IMR $\mathrm{Kerr_{221}}$ $\mathrm{EW1_{221}}$ $\mathrm{EW1_{221}}$
GW150914 $68.8^{+3.6}_{-3.1}$ $71.7^{+13.2}_{-12.5}$ $65.2^{+10.1}_{-8.0}$ $0.69^{+0.05}_{-0.04}$ $0.69^{+0.18}_{-0.36}$ $0.55^{+0.20}_{-0.25}$ $\leq 0.65$
GW190521 $256.6^{+36.6}_{-30.4}$ $284.0^{+40.4}_{-43.9}$ $284.7^{+34.4}_{-28.4}$ $0.71^{+0.12}_{-0.16}$ $0.78^{+0.10}_{-0.22}$ $0.78^{+0.09}_{-0.12}$ $\leq 0.36$
GW190521_074359 $88.1^{+4.3}_{-4.9}$ $86.4^{+14.1}_{-14.8}$ $84.7^{+11.7}_{-9.9}$ $0.72^{+0.05}_{-0.07}$ $0.67^{+0.17}_{-0.34}$ $0.63^{+0.16}_{-0.23}$ $\leq 0.58$
GW200224_222234 $90.3^{+6.4}_{-6.3}$ $88.6^{+15.5}_{-15.2}$ $84.0^{+12.4}_{-9.3}$ $0.73^{+0.06}_{-0.07}$ $0.60^{+0.23}_{-0.42}$ $0.50^{+0.23}_{-0.25}$ $\leq 0.70$
Figure 6: Posterior distribution for the remnant mass and spin resulting from the analysis of GW150914 for three different ringdown templates and the results from the full LVK IMR model using IMRPhenomXPHM (obtained from the data release [ligo_scientific_collaboration_and_virgo_2022_6513631]).
Figure 7: Posterior distributions for the remnant spin $\tilde{a}_{f}$ and magnetic field $\tilde{B}_{f}$ resulting from the anlalysis keeping the remnant mass fixed to the IMR median value: $M_{f}=68.8M_{\odot}$. Due to the restriction on the remnant mass, there is enhanced sensitivity to $\tilde{B}_{f}$.
Having described the essential tools of our methodology, we now shift to analyzing LVK data provided by the GW Open Science Center [KAGRA:2023pio, LIGOScientific:2019lzm]. Employing the described framework, and assuming that the remnant object to be a EW1 remnant described by scalar quasinormal modes, we conduct an analysis of the ringdown data for four binary black hole merger events from the LVK catalogue. The events were chosen based on their high signal to noise and sufficiently long waveform for a ringdown-only analysis.
We initially adopt simple uniform prior distributions for the remnant mass, spin and magnetic field, $M_{f}\in[10,500]M_{\odot}$, $\tilde{a}_{f}\in[0,0.99]$, and $\tilde{B}_{f}\in[0,1]$ respectively. Note that we are choosing to extrapolate the linearized analysis to form the EW1 template over the full range of spin and magnetic field. It is clear that the extrapolation has limited applicability as one approaches large values of $a$ and $B$, but this will be sufficient to explore sensitvity to $B$ as a nuisance parameter. Finally, we also account for the choice q=2aMBq=-2aMB to ensure overall electrical neutrality, we address the extremality bound $q^{2}+a^{2}\leq M^{2}$ by imposing an a priori joint limit on the spin and magnetic field parameters $\tilde{a}^{2}_{f}+4\tilde{a}_{f}^{2}\tilde{B}_{f}^{2}<0.99$, excluding near-extremal black hole configurations.
We also consider a more constraining choice of priors, referred to as the IMR prior [Laghi:2020rgl], with uniform distributions of $M_{f}$ and $\tilde{a}_{f}$ containing the $99\%$ credible intervals of their one-dimensional probability density functions determined from a full LVK IMR analysis of the event. For GW150914 this set of priors corresponds to: $M_{f}\in[62,75]M_{\odot}$ and $\tilde{a}_{f}\in[0.55,0.75]$. This choice assumes that the nuisance parameter $\tilde{B}$ is weakly correlated with other black hole parameters, and that the IMR gravitational wave emission follows the Kerr template. This approach naturally provides a more stringent constraint on the value of $\tilde{B}_{f}$.
The results are shown in Figs. 5, 6 and 7. The corner plots show in Figs. 5 exhibit the $90\%$ credible interval of the two-dimensional posterior distributions on remnant redshifted mass, spin and magnetic field for the set of four selected events using ths simple uniform priors. The results of which are also tabulated in Table 4, where they are compared to results determined by LVK using the standard Kerr template. Note that care is needed in comparing the inferences from different templates. Kerr221 and EW1221 apply just to the ringdown signal from pyRing, while IMR is the full Kerr template applied also to the inspiral and merger. The inferred central values of mass and spin therefore differ, but the results for Kerr221 and EW1221 agree with each other and with IMR within their respective 90% confidence intervals.
@@
It is important to emphasize that, due to the extrapolation of the linearized EW1 template, the actual constraints on $\tilde{B}$ are properly understood as indicative of the level of sensitivity to $\tilde{B}$ as a nuisance parameter, and not a true constraint on the magnetosphere. The results nonetheless indicate no evidence for a deviation from the Kerr hypothesis, and while the ensuing bounds are relatively weak, the current data already provides sensitivity to magnetic fields above $\tilde{B}\sim 0.3$. This is still very large in astrophysical terms, but indicates that the presence of a magnetosphere may provide one further factor complicating the extraction of black hole parameters from the merger. The exploratory nature of this analysis, apparent in the approximations A1 and A2, limits the scope of broader conclusions, but relaxing both assumptions is technically feasible, as discussed in the next section.
@@ 5 Concluding Remarks
[…]
@@ Acknowledgements.
We are particularly grateful to Gregorio Carullo for assistance with the implementation of pyRing, and to Michel Lefebvre for many valuable discussions regarding aspects of this work and for carefully reading the manuscript. We would also like to thank Emanuele Berti for illuminating discussions at the Testing Gravity 2023 conference, Frans Pretorius for providing additional insight at the Werner Israel Memorial Symposium, and Michalis Agathos for helpful comments at the poster session at the Kavli-Villum Summer School on Gravitational Waves. This work was supported in part by NSERC, Canada, and has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/), a service of the LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. Software: This work makes use of the following python packages: pyRing, corner, cpnest, GWpy, lalsuite, matplotlib, numpy, scipy, and seaborn [pyRing, cpnest, gwpy, lalsuite, matplotlib, numpy, scipy, seaborn].
@@ Appendix A Ernst-Wild asymptotics at ${\cal O}(\tilde{B}^{2})$
and the Klein Gordon equation becomes [Brito:2014super],
@@
The effective potential for the Schwarzschild black hole on the first line is only corrected by a constant at large radius at ${\cal O}(\tilde{B}^{2})$. The shift is exhibited in Fig. 8, with the shaded region showing the domain beyond the Melvin radius where the asymptotics would be modified more substantially at ${\cal O}(\tilde{B}^{4})$. The restriction to ${\cal O}(\tilde{B}^{2})$ thus restores the expectation within the Kerr spacetime that the QNM spectrum is primarily affected by the behaviour of the potential near the horizon and photon sphere of the black hole [Konoplya:2006ar, Konoplya:2006rv, Konoplya:2007yy, Konoplya:2008hj, Brito:2014super].